Micromedia offers computer training seminars on a variety of topics. In the seminars each student works at a personal computer, practicing the particular activity that the instructor is presenting. Micromedia is currently planning a two-day seminar on the use of Microsoft Excel in statistical analysis. The projected fee for the seminar is $650 per student. The cost for the conference room, instructor compensation, lab assistants, and promotion is $9300 per 2 days. Micromedia rents computers for its seminars at a cost of $150 per computer per day.

  1. Develop a model for the total cost to put on the seminar. Let x represent the number of students who enroll in the seminar. Round your answers to the nearest whole number. For subtractive or negative number use a minus sign even if there is a + sign before the blank. (Example: -300)

    Total cost =  + ·x
  2. Develop a model for the total profit if x students enroll in the seminar. Round your answers to the nearest whole number. For subtractive or negative number use a minus sign even if there is a + sign before the blank. (Example: -300)

    Total profit = ·x +
  3. Micromedia has forecasted an enrollment of 40 students for the seminar. How much profit will be earned if their forecast is accurate? Round your answer to the nearest dollar.

    Total profit = $  
  4. Compute the breakeven point. Round your answer to the nearest whole number.

    The breakeven point is  students.
Answer #1

(a)

Total cost = 9300 + 150*2.x = 9300 + 300.x

[Note that $150 is the per day cost and the program is for 2 days, so on a per-student basis, this cost is $300]

(b)

Total profit = Total revenue - Total cost = 650.x - (9300 + 300.x) = 350.x + (-9300)

(c)

For x = 40,

Total profit = 350*40 + (-9300) = $4,700

(d)

At the breakeven, the total profit = 0

So,

350.x + (-9300) = 0

or, x = 9300/350 = 26.57 or 27 students